CS294-3: Algorithms in the real world (Fall 97) Course questionnaire

For each of the following questions answer with a number between 1 and 5.

1. The lectures are to slow (1), or to fast (5).

2. Would you rather see more detail on individual topics with fewer topics (1), the current level of detail (3), or less detail with more topics (5).

3. The lecture notes are not helpfull (1), or very helpfull (5).

4. The readings are not helpfull (1), or very helpfull (5).

5. The web pages are not helpfull (1), or very helpfull (5).

6. The assignments are too easy (1), or too hard (5).

7. For the purpose of helping to understand the material the assignments are not helpfull (1) or very helpfull (5).

8. The criteria for grading is unclear (1), or clear (5).

9. The class is badly organized (1) or well organized (5).

Please add any other comments you have in the space below. Feel free to mention topics you would like to see covered (although I can’t make any promises).